OpenStudy (anonymous):

What is the simplified form of (y/(y^2-x^2))-(x/(y^2-x^2))

OpenStudy (anonymous):

\[\frac{y}{y^2 - x^2} - \frac{x}{y^2 - x^2}\]

OpenStudy (anonymous):

ya that

OpenStudy (anonymous):

Do you know about the difference of two squares?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Factoring the difference of two squares states that: \[a^2 - b^2 = (a + b)(a - b)\] Can you see any matching patterns with your expression?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

is the answer 1/y-x?

OpenStudy (anonymous):

\[\frac{y - x}{y^2 - x^2} = \frac{y - x}{(y+x)(y - x)}\] Check your cancellation more carefully

OpenStudy (anonymous):

1/y+x?

OpenStudy (anonymous):

Good job